OpenStudy (anonymous):

Can someone help me walk through this problem? Solve the system by substitution. -x -y -z= -8 -4x +4y +5z =7 2x +2z =4

OpenStudy (anonymous):

use the 3rd option, and make it 2x= 4-2z--> x=2-1z

OpenStudy (anonymous):

substitute into option1, then you will have option 1 comprised of y and z. Make isolate for y (so it looks like y=~~~z) and substitute that all into second equation. x must be in terms of z, y must in terms of z, and z is left as it is. Then you have ~~z= 7 so you get z, then you can get values of other two variables as you know ~y= ~~z and ~z= ~~x

OpenStudy (anonymous):

Thank you so much!